OpenStudy (anonymous):

Rewrite 3x + 4y = 10 using function notation

OpenStudy (raffle_snaffle):

Solve for y and y=f(x)

OpenStudy (texaschic101):

solve for y and what do you get ?

OpenStudy (texaschic101):

3x + 4y = 10 --- subtract 3x from both sides 4y = -3x + 10 --- divide both sides by 4 y = -3/4x + 10/4 y = -3/4x + 5/2 now replace y with f(x) f(x) = -3/4x + 5/2 ==> function notation

OpenStudy (anonymous):

thnx